HUBBLE SPACE TELESCOPE DAILY REPORT #2384 PERIOD COVERED: 0000Z (UTC) 06/01/99 - 0000Z (UTC) 06/02/99 Daily Status Report as of 153/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7949 (CCD Dark and Bias Monitor -- Cont. to June 99) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a CCD dark calibration and a bias monitoring program. It was also scheduled to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Seven Sets of STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.3 Completed Three Sets of WF/PC-2 8053 (WF/PC-2 Cycle-7 Earth Flats -- Part II) The WF/PC-2 was used to take Earth flat calibration images. This proposal obtains sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed S/C 8536 (GS Acq test for 8267) A guide star acquisition was conducted using Target Star #1. The is a robust test of the guide stars to be used by proposal 8267. The acquisition was executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 7379 (The Dense Core of M-30) The WF/PC-2 was used to study the dense post core collapse globular cluster NGC-7099 (a.k.a., M-30). These studies will be combined with our ground-based observations of M-30 to help formulate a coherent understanding of this remarkable cluster, teaching us more about the dynamics of dense evolved clusters and the nature of the exotic objects once hidden in their cores. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/CCD/MA1/MA2 7757 (Time-Resolved Spectrophotometry of the Smallest Mass Function Binary Star) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make time resolved spectrographic observations of the x-ray, small mass, contact binary star KZ-TrA (a.k.a., The 7s Pulsar). The optical counterpart of this star exhibits optical pulsations with the same period as the X-ray pulsations. As presented in HSTAR 7153 and 2.1, there was a loss of lock during this proposal, and the take data flag was down temporarily. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 6 As documented in HSTAR 7153, there was a loss of lock at 152/222610Z. Although lock was re-acquuired on the first attempt, the STIS bail-out flag was issued. The proposal described in 1.7 was affected. The acquisition at 153/045141Z failed to RGA control when the search radius limit was exceeded on both FGS-2 and FGS-3. The preceding full maneuver update was good and the subsequent FHST map showed good attitude. HSTAR 7155 wqas written. Three proposals, STIS 7949, FGS 8334 and WF/PC-2 8090, occurring during the period of the next report, were affected. Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared four times. At 152/222700Z, the FGS-2 motor encoder temperature was flagged out of limits (high). The upper limit is 23.0 degC and the temperature briefly registered 23.15 degC. HSTAR 7154 was written. Per an operations request, the SSR engineering storage boundaries were increased at 153/0205Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH